Greenhab Report – January 29th

Prepared by: Julia DeMarines

Environmental control: Heating

Ambient with door opening: Shade cloth on

Working Hour: 19:00
Inside temp at working hour: 18° C
Outside temp during working hours: 1° C
Inside temperature high: 34 C
Inside temperature low: 16 C
Inside humidity: 23 %RH
Inside humidity high: 46 %RH
Inside humidity low: 15 %RH

Hours of supplemental light:
For the crops 05:00 to 11:59 PM

Changes to crops: moved some of the tomato plants and bean plants to the blue aquaponics container.

Daily water usage for crops: 8 gallons

Time(s) of watering for crops: 19:00

Changes to research plants: Need more info about existing research plants.

Aquaponics: Crew 188 will not be attempting to assemble the aquaponics project.

Narrative: This is Julia’s first day as Green Hab officer. She watered the plants and removed some of them from the wooden platform and put them in the blue aquaponics container after recommendations from Shannon. She went through changes to the plants that Shannon would like to have done to the Green Hab and thought about a plan moving forward. Tomorrow she will begin implementing the plan and she looks forward to some redecorating.

The temperature sensor is still currently malfunctioning. The tens digit has missing sections. Julia would like to know if there is a certain time(s) that the temperature and humidity need to be measured daily and specific information on ongoing projects that Julia needs to maintain or monitor.

Sol Summary – January 29th

Sol 1

Summary Title: Kicking up Martian Dust
Author’s name: Ryan L. Kobrick, Ph.D., MDRS Crew 188 Commander

Mission Status: All nominal

Sol Activity Summary:

The mission kicked off 1:01:01 local time after a morning of spacesuit training and navigational discussions of the exploration zone. As the terrain changes, each subsequent crew on the "red planet" adjusts to imposed closures of some roads, while other pathways open for exploration. The crew snapped team photos with four nations and four flags, (Canada, USA, Australia, and Japan), and with the one uniting Mars Red-Green-Blue flag.

The crew planned and executed two EVAs geared towards a technical shakedown of equipment and MDRS traditions located at "Marble Ritual". Three crewmembers on each EVA drove surface vehicles to a predetermined destination to make sure the surface suits were adjusted, vehicles driven suited (the ATVs and Rovers), and then on-foot exploration. Additionally, cameras, GPS units, biomedical devices, tablets, and specialized gloves were all tested in Martian analog conditions.

The relived to be underway the crew also realizes that the clock is ticking on their mission and there is a lot of data to collect and share. Things are just getting ramped up!

Spacesuit Up!
Ryan L. Kobrick, Ph.D.
MDRS Crew 188 Commander

Look Ahead Plan:
In the morning three crew members will explore the main superhighways to the North of hab. This will help ground truth critical waypoints, calibrate GPS equipment, familiarize the crew with the regions, and ensure safe exploration of all future EVAs. Crewmembers will be working on calibrating the telescope and 360 camera for the upcoming astronomical mega event of the lunar eclipse of the January Blue Moon.

Anomalies in work:
MDRS Director is coordinating power system updates with contractors.
Robotic observatory currently not functional. MDRS Astronomy lead working problem.

Weather:
A warm day on "Mars" with crewmembers comfortable on EVA.

Crew Physical Status:
The crew is learning how to combat dehydration in this new environment.

EVA:
Two EVAs completed with three crew on each.

Reports to file:
1. Sol Summary
2. Operations Report
3. Journalist Report
4. HSO Checklist
5. EVA Request for Sol 2
6. Daily Photos
7. Mission Summary
8. EVA 001 and EVA 002 Report
9. Green Hab Report

Support Requested:
None.

EVA Report – January 29th

EVA #2

EVA Date: 29 Jan 2018

Crew Executive Officer Reporting

Purpose of EVA: Spacesuit activity recognition and assessment of in-suit performance
Location of EVA: South of Marble Ritual (walking), Pooh’s corner (vehicle)
UDM27 Coordinates: 518800 E, 4250600 N

Number of EVA Crew: 3

Crew Members going on EVA:
Renee Garifi (Executive Officer), Julia De Marines (Crew Astronomer), Zac Trolley (Crew Engineer)

Rovers used:
– DEIMOS
– Opportunity
– Spirit

Rover time usage: 0.5 hours

Duration of EVA: 1.3 hours

EVA departure time: 15:10
EVA return time: 16:42

Time Checks:
Start Prep: 15:32
Airlock Start: 15:10
Roll out: 15:20
Check-In: 15:32
Comms on Overshoot: 15:40
Scrub due to helmet issue: 15:54
Renee in Airlock: 15:58
Zac and Jules return to get Rover: 15:59
Jules in Airlock: 16:15
Zac in Airlock 16:28
Equipment Cleaned and put away: 4:42

Summary: The Crew attempted to locate the Marble Ceremony landmark near the Pooh’s Corner rock area. They attempted to utilize one of the hab GPS units to identify the planned EVA coordinates. We took with us one hab GPS unit, as well as one of the small, laminated MDRS, printed maps to find our waypoints. Our objective was to perform a training EVA to familiarize our crew (in two team groups of three) with EVA operations first-hand from a participant perspective. All members brought cameras for photos and video footage to send to Mission Support. All three rovers used were fully functional and tested in the rough paths of the Martian Terrain on Cow Dung Road.

We experienced an off-nominal situation with one crewmember during this second EVA of the day. All three of the crew experienced a higher-than-normal amount of helmet fogging while walking. It was realized shortly after parking the rovers on cow dung road that we had overshot the EVA destination and were too far north. While walking back to the rovers, one crewmember experienced helmet fog to the point of having impaired visibility and was unable to drive their rover. After waiting a few minutes for the fogging to clear, the crewmember began to have sweat drip down into the eyes and collect in the glasses. They requested assistance from another crewmember to turn the rover around for them. In the process, the crewmember having the fogging issue experienced a total visual loss when their hat fell from their head and into the front of their helmet, completely blocking their view.

When this hat anomaly occurred, the crewmember made the call to be driven back to the hab to receive assistance with the helmet from the HabCom team. All three crewmembers returned on two rovers, dropping the impaired crewmember at the airlock and returning to retrieve the third rover.

The remaining two crewmembers received permission to resume the EVA and made it out to Pooh’s corner safely and completed the marble ceremony. Despite having to return one crewmember to the hab early, this EVA was highly successful for three reasons. First, it provided a non-life threatening emergency that required immediate assistance for the crew to respond to.

There is a long list of reasons why an EVA will terminate early, so we are fortunate that this instance was minor and could be resolved by removing the helmet back at the hab. Secondly, the crew of three proved they could work together in an off-nominal situation to problem solve and quickly make a decision that places the safety of the crew above the objective of the EVA. Lastly, it allowed us to demonstrate our contingency plan for returning to the hab if one of the rovers breaks down since one rover had to remain parked while all crew returned to the hab.

The crew engineer extended their EVA in order to service the ATVs and Rovers that we used during the day.

EVA Report – January 29th

EVA #1

EVA Date: 29Jan2018

Health & Safety Officer Report

Purpose of EVA: Development of Situational Awareness for Spacesuit Activity

Location of EVA: South of Marble Ritual (walking), Pooh’s corner (vehicle)

UDM27 Coordinates: 518800 E, 4250600 N

Number of EVA Crew: 3

Crew Members Going on EVA: Ryan Kobrick (Commander), Sarah Jane Pell (Residence-in-Artist), Tatsunari Tomiyama (Health & Safety Officer)

ATV used: 1,2,3

ATV usage: 5 minutes

EVA Duration: 1 hour 7 minutes

EVA Departure Time: 13:30

EVA Return Time: 14:32

Time Check:

13:01 Prep Start + Simulation Start
13:25 Airlock
13:30 EVA Start
13:35 ATV Leave
13:40 Parked
14:02 Arrived Ritual Marbles
14:18 Return to Hub
14:27 Arrived Parking Lot
14:32 Airlock

Summary: Crews attempted EVA to develop situational awareness for space suit performance. We used 3 ATVs to travel to the targeting location; Marble Rituals. When we arrived at the nearest place on Pooh’s corner, we parked ATV and started to walk. We found a fossil and recent animal footprint soon. Once arrived, we take photos and investigated the location. Then, we returned to the Hub. Overall, there was no significant issue to this EVA and we enjoyed the first simulation.

Operations Report – January 29th

Prepared by: Zac Trolley, Operations Manager

Non-nominal systems: All systems are operating in their handover condition
Notes on non-nominal systems: None.

Generator (hours run): 17.1h
Generator turned off: 10:23
Generator turned on: 17:38

Solar
SOC 7:37 93%
SOC 17:38 85%

Diesel: 83%
Propane: Information to be provided by the Director
Ethanol Free Gasoline (5 Gallon containers for ATV): 8 Gallons
Water (trailer): 0 Gallons
Water (static): 400 Gallons
Trailer to Static Pump used: No
Water (loft) – Static to Loft Pump used: Yes
Water Meter: @17:00 130497.5
Toilet tank emptied: No
ATVs Used: HONDA, 350(1), 350(2), 350(3)
All ATV’s used for a short EVA to the marbles. No fuel was added.
Oil Added: added to ATV #2
ATV Fuel Used: None
# Hours the ATVs were used today: 0.5
Notes on ATVs: None

Deimos rover used: Yes
Hours: 106.9h
Beginning charge: 100%
Ending charge: 100%
Currently charging: no

Spirit rover used: Yes
Hours: 16.4h
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Opportunity rover used: Yes
Hours: 13.5h
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Curiosity rover used: Yes
Hours: 7.9h
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

HabCar used and why, where? No
General notes and comments: None
Summary of internet: All nominal
Summary of suits and radios: We had an issue where one of the Radios got stuck on transmit on channel 12, causing a comms blackout. This happened with everyone in the Hab, getting ready for an EVA. The Radio was found and the problem was solved before the EVA or anyone left the Hab.
Summary of Hab operations: All nominal
Summary of EVA operations: Had an issue with the facemasks fogging up. Several people had issues with visibility, one person needed to be driven back due to visibility. We will investigate an airflow solution to provide more anti fogging ability.
Summary of GreenHab operations: All nominal
Summary of Science Dome operations: All nominal.
Summary of RAM operations: Not Operational.
Summary of health and safety issues: All nominal

Questions, concerns and requests to Mission Support: Quality data comes from the consistent method of collection of data. To that end, I have a few questions about the data being collected.

For the state of charge, what is that data being used for, and when should it be taken? Without the SOC being tied to the time of day and amount of sun, it doesn’t have much value. (unless I’m missing something)

The water number has a similar issue. Knowing what time of day it was taken, and taking it at the same time every day provides more useful data. I’ve been writing down the water flow several times a day to track our own use. What would you like the information to accomplish?

Last, the information on the rovers won’t be accurate, meaning if I plugged them in, the battery percentage won’t be accurate at the time of this report. Is that to keep an eye on the battery function? If so, I will only transmit information that is relevant to that investigation.

Sol Summary – January 28th

Crew 188 Sol 0 Summary Report 28JAN2018

Sol 0

Summary Title: Training for Mars
Author’s name: Ryan L. Kobrick, Ph.D., MDRS Crew 188 Commander
Mission Status: All nominal

Sol Activity Summary:
Today was a packed day starting with the early departure of Crew 187. The crew moved into the state rooms and conducted another water refill in the morning to Hanksville. Training with MDRS Director Shannon Rupert covered MDRS rules, philosophy, and key systems. The crew conducted an EVA test of the ATVs and Rovers in preparation for our simulation. The crew united by a common passion for sharing their experience has started sharing highlights with the common hashtag #MDRS188.

Spacesuit Up!
Ryan L. Kobrick, Ph.D.
MDRS Crew 188 Commander

Look Ahead Plan:
The crew will complete training tomorrow morning with the EVA spacesuit systems and general navigation of the local area. The "landing" and simulation will kick-off at approximately High Noon local time. The crew is planning two short EVAs with half the crew (3) on each. Each EVA will use hybrid driving/walking modes to familiarize the crew with nominal operations with the surface spacesuits. Reports and scientific projects will be prepared in the afternoon.

Anomalies in work:
MDRS Director is coordinating power system updates with contractors.
Robotic observatory currently not functional. MDRS Astronomy lead working problem.

Weather:
A beautiful day. Cool and clear.

Crew Physical Status:
Some dry air dehydration symptoms. See HSO Report.

EVA:
ATV/Rover training with Shannon.

Reports to file:
1. Sol Summary
2. Operations Report
3. Journalist Report
4. HSO Report
5. EVA Request
6. Daily Photos

Support Requested:
Creative request information will be coming.

Operations Report – January 28th

Operational Report

Crew 188: Zac Trolley

28/01/2018

Non-nominal systems: None

Notes on non-nominal systems:

Generator (hours run): 14h

Generator turned off: 9:50

Generator turned on: 17:45

Solar— SOC

@9:50 80%

@13:00 100%

@17:45 82%

Diesel: 85%

Propane: 75%

Ethanol Free Gasoline (5 Gallon containers for ATV): 10 Gallons

Water (trailer): 0 Gallons

Water (static): 400 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Water Meter: @19:00 130469.3

Toilet tank emptied: No

ATVs Used: HONDA, 350(1), 350(2), 350(3)

Oil Added? No

ATV Fuel Used: Very little in learning how to fuel the ATVs

# Hours the ATVs were used today: 0.5h Training

Notes on ATVs: Took the ATVs on a training run. The hours run have not been updated from yesterday.

Deimos rover used:

Hours: 102.3 (crew) + Director discretional hours

Beginning charge: 100%

Ending charge: –

Currently charging: Maybe

Sojourner rover used: Assigned to director only.

Hours: Director discretional hours

Beginning charge:

Ending charge: –

Currently charging: Maybe

Spirit rover used: Yes

Hours: 15.1h

Beginning charge: 100%

Ending charge: –

Currently charging: Yes

Opportunity rover used: Yes

Hours: 12.4h

Beginning charge: 100%

Ending charge:

Currently charging: No

Curiosity rover used: Yes

Hours: 6.6h

Beginning charge: 100%

Ending charge: –

Currently charging: Yes

HabCar used and why, where? Yes, Water run.

General notes and comments: There is a faint smell of propane in the Hab, we believe it may be from the heater on the lower floor when we tried to turn it on. There was also a minor smell of fuel gas during the day.

Summary of internet: All nominal

Summary of suits and radios: Not checked

Summary of Hab operations: All Nominal

Summary of EVA operations: N/A

Summary of GreenHab operations: All nominal

Summary of Science Dome operations: All nominal.

Summary of RAM operations: Not Operational.

Summary of health and safety issues: All nominal

Questions, concerns and requests to Mission Support: N/A

Journalist Report – January 27th

[draft status]

English

[Sol 15]

[A new Hope]

People say that good is for a short time. Two weeks, 15 sol have passed and now we are in the opposite place. The 188 crew has arrived on this planet. You will see them smile, sigh and miss as has happened with us.

But to get to this point of the day it is necessary to count the cleaning routine of today. 8 am, brooms and mops ready. Greenhab, Science Dome, Observatory and Hab ready. The suits and helmets in place, clean kitchen and suitcases made.

The International Space University team arrived and our work as hosts began. It’s time to show them the site, tips, ideas and warnings. When we finished we decided it was time for our first meal on Earth and their last meal before Mars in some peculiar Utah-style restaurant in the middle of a Martian road. Milkshakes, hamburgers, fries and chicken strips, the truth maybe is that we only want to recover the lost weight.

Back home the last night begins and the only thing left is to enjoy these hours. I write these final words surrounded by my favorites, planning new projects and collaborations. We do not want to lose ourselves back to Earth.

And now personal feelings but that are shared here as oxygen. This last part is dedicated to Attila, Cynthia, Danton, David, Luis and Oscar, my dear astronauts, my dear Martians.

Thank you for being like that, for laughing and enjoying this time. For being young but one of the most mature and intelligent people that you could know, for being sensitive, educated and so honest, for taking care of us and our humor along with the chocolates, for your madness, fun and friendship, for your nobility, solidarity and companionship, for your intelligence, character and food.

Thank you for allowing yourself to feel and open your minds and hearts to others. We will miss ourselves along with our defects, virtues, personalities and emotions.

Sometimes it happens that from those who least expect something you get the best results. Congratulations astronauts we achieve the goal, we fulfill the mission and now there is nothing left to do here. Soon we will return together or separately, in the near of distant future. Goodbye Mars, hello Earth!

LATAM II thanks all its collaborators and see you soon.

Tania Robles, MDRS Crew 187 Journalist

Sol Summary – January 27th

Crew 187 Sol 15 Summary Report 27JAN2018

Sol 15

Summary Title: Beyond the Terra

Author’s name: Cynthia Fuertes Panizo

Mission Status: All nominal

Sol Activity Summary:

Hello Earth,

How much has the Earth changed since the Sol 1?

It’s hard to explain in words all the feelings. Sadness and happiness at the same time. Let’s explain to you what we did during our last day, then there will be time for farewells. Today we woke up at 07:30, we ate our last breakfast here on Mars, and started to clean all the Hab, the lower and the upper deck, the GreenHab, the science dome and the observatory. Every Martian was running around cleaning everything, keeping their things, cooking and so on. Then Dr. Ruppert came back to check everything. At 15:00, Atila, Danton and I went to welcome the Crew 188 from ISU… Welcome guys to Mars! We hope you really enjoy living here as much we did! When every Martian arrives at the Hab we did a tour for Crew 188 showing them the GreenHab, the science dome, the Musk observatory and everything here, our home. Later, Shannon visited us to explain to crew 188 how is living on Mars. Meanwhile, Danton flew his drone around the Hab to do the aerial 3D mapping. And, at 17:00, we teleport to Earth for eating hamburgers and chicken… Yummy! A long time ago that we didn’t eat terrestrial food. When we came back to Hab, we started to write the reports for capCom and we answered doubts that could have the crew 188. By the end of the night, crew 187 are together talking about our projects, our lives, and so on … as a Martian Family!

It’s time to say goodbye! This place was our home, our sweet home. We have a lot of good memories here on Mars. I was lucky to have as my crew, the best Martians in the Universe. It was an honor to be the commander of the crew 187. Every Martian was a key to survive to this great adventure that is to live on Mars. Atila, Luis, David, Oscar, Danton, and Tania… No matter wherever I stay, on Earth or on Mars, I will always love you. You are my Martian family! We are going to be now and forever: CREW 187 – LATAM II !!!

Ad Astra

Cynthia Fuertes Panizo

Commander Crew 187 – MDRS

Look Ahead Plan:

Arrive at Earth! See other people. And continue working on space issues.

Anomalies in work: None

Weather: Temperature: -3°C, Fair-weather, Winds speed: Calm, Humidity 36%, Barometer 30.43 in

Crew Physical Status: Healthy

Reports to be filed: Sol Summary, GeenHab Report, Operations Report and Journalist Report.

Support Requested: None